Dear Richard dear all,

please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO were:

OAA: Armando, Marco, Runa
LBTO: Guido, Richard Sosa
MG: Biasi

Total # of person: 7

W unit

no activity

AO control SW

no update


Thin Shell
1) calibrated and tested reproducibility of the centering tool for the shell to be used on the flipping machine
2) shell support has been moved from shipping frame to flipping machine and centered on it

Secondary Unit
1) test of capacitive sensors:
- 5 external actuators have been found not working:
- 3 actuators were fixed changing the signal ribbon cable of one of them
- 2 were found having the elastic contact picking the capacitive sensor signal in contact with the actuator body (ground). The problem was solved bending the elastic contact. We checked that the bending operation did not change the capacitive sensor reading (difference within 1 bit,~3nm).
- the problem with the touching of the elastic contact showed a reduction of the clearance between the actuators and the reference body, possibly due to an increase of the clocking error between cold plate and reference body (Creeping? Transportation?). A group of actuators around the faulty ones where very critical, highly possible they would touching using the unit in the cold. The elastic contact has been bent also for them always checking no significative variation of calibration.

Retroreflector installation
- Discussed with John Little the procedure for installing the retroreflector holder on the telescope


The other people from Arcetri did not go up to the mountain because of deep snow on the road.

Next update tomorrow.

All the best


-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback